Two times a number plus three times another number equals 13. The sum of the two numbers is 7. What are the numbers ?
Answer by praseenakos@yahoo.com(507)

sum of numbers = 7
if one number is 'n'
other number is 7-n
given, Two times a number plus three times another number equals 13
==> 2n + 3(7-n) = 13
==> 2n + 21 -3n = 13
==> 2n - 3n + 21 = 13
==> 1n + 21 = 13
n + 21 - 21 = 13 - 21
==> n = -8
one number is -8
so the other number is 7 - (-8) = 15